two:balls of:mass m1 and m2 are separated from each other by a powder charge placed between them The whole:sysytem:is at rest on the ground suddenly the powder charge is explodes and masses are pushed apart The:mass m1 traveles a dostance s1 and stops If the coefficient of friction betweenthe balls and ground is:same the mass m2 stops:after travelling

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Let the velocities of the balls just after explostion bev_1,v_2.
Hence from conservation of linear momentum,

0=m_1v_1+m_2v_2
\implies v_2=-\dfrac{m_1}{m_2}v_1

The deceleration of both the balls due to friction=a=-\mu g

Hence writing the equations of motion of the balls gives,

0^2-v_1^2=2as_1\implies s_1=\dfrac{v_1^2}{2\mu g}
for the second ball,

s=\dfrac{v_2^2}{2\mu g}=(\dfrac{m_1^2}{m_2^2})s_1